bella2025:

y=14x+4 y= -2x-5 -4,3 x=-4 y=3

bella2025:

would my answer be correct @axie

axie:

One sec, graphing it looked weird so im going to try plugging in your answers, afk for like 2 mins

axie:

@extrinix he can help, i have class now,

bella2025:

@axie wrote:
@extrinix he can help, i have class now,
ok thanks

Extrinix:

What is the question?

bella2025:

y=14x+4 y= -2x-5 -4,3 x=-4 y=3 I wanted to know if I got it right

Extrinix:

That's equations, not a question, I'm confused by that

bella2025:

@extrinix wrote:
That's equations, not a question, I'm confused by that
did I do the equation right

Extrinix:

@bella2025 wrote:
@extrinix wrote:
That's equations, not a question, I'm confused by that
did I do the equation right
There's 2 equations, \(y=14x+4\) and \(y=-2x-5\) Are you trying to find the point at which they meet?

bella2025:

@extrinix wrote:
@bella2025 wrote:
@extrinix wrote:
That's equations, not a question, I'm confused by that
did I do the equation right
There's 2 equations, \(y=14x+4\) and \(y=-2x-5\) Are you trying to find the point at which they meet?
I did the equations in my calculator online and i got that they meet at -4,3 but im not sure its right

Extrinix:

Assuming that you're trying to find the points at which the lines meet, you would be correct

bella2025:

@extrinix wrote:
Assuming that you're trying to find the points at which the lines meet, you would be correct
thank you I got more questions to